OpenStudy (anonymous):

Algebra help please!! 9. Which of the following points are not solutions of 2y-5x>6? I. (-3,4) II. (5, - 5/3) III. (3/4, 5) IV.(2.6, 5) A. I and III B. III only C II and IV D. III and IV

OpenStudy (usukidoll):

we are given a pair of points in the form of (x,y) so for I. we have (-3,4) so x = -3 and y = 4 plug that into the equation and see what happens. Does it satisfy the inequality problem of 2y-5x > 6 or not?

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

Is C correct??

OpenStudy (anonymous):

Find the slope of each line. A. A line that is parallel to the graph of Y=1/2 X +7. B. A line that is perpendicular to the graph of Y= -2x -3.

OpenStudy (anonymous):

Can I have help with that one too please :)

OpenStudy (usukidoll):

I need to check on whether or not C is correct as I am eating lunch at the moment. Also for the second question (which by the way you should open a new question) an example of 2 perpendicular lines is  so you need to graph y = -2x-3 and find another equation that will intersect with y = -2x-3 parallel lines look like this  there is no intersection so graph y = 1/2x+7 first and then find another equation that's similar to y =1/2x+7